OpenStudy (anonymous):

Find the height of a rectangular prism if the surface area is 458.17 in2, the width is 6.2 in and the length is 8.3 in.

OpenStudy (nurali):

Let: SA = Surface Area = 458.17 in² L = Length = 8.3 in W = Width = 6.2 in H = Height = What we're solving for. The formula for the surface area of a rectangular prism is SA = 2*(LW + LH + WH) Substituting values: 458.17 = 2*(8.3*6.2 + 8.3H + 6.2H) 229.085 = 51.46 + 14.5H 14.5H = 229.085 - 51.46 H = 117.625/14.5 H = 12.25 inches.